Test of Hg loop and solenoid @ MIT
26. Feb. 2007 (Monday):
Charging the HPU with hydraulic fluid
27. Feb. 2007 (Tuesday):
Lowering the Hg loop into the pit: Due to limited space the lowering of the Hg loop was tricky. The Hg loop is placed with the snout half way into the bore (video). Rigging of the mercury loop is possible via two rectangular guides welded to the Hg box.
The HPU is installed at the top of the pit.
During the following days the solenoid is permanently cooled to 80 K (fluctuating up to about 110 K). The bore is actively heated by the means of a bore heater (approx. 200 W max). Without the heater the bore accumulates ice.
28.Feb. 2007 (Wednesday), AF arrived
Opening the top cover in order to allow installation of fiber optics (video). Before hand the secondary volume is flushed actively. The mercury level is always zero at any time (during the whole tests at MIT, signal only shown upon exposure to open mercury flasks; an offset of about -48 micro-gram/m3 is occuring).
An overview picture of the installation before preparing to fill the mercury.
Connecting the 480 V supply for the HPU.
Last issue in the morning was to pull the hydraulic lines from the HPU to the mercury loop (video: connecting the hydraulic hose). Special attention has to be given to the interference of the hydraulic lines with the fiber optics, which are very fragile. The installation of these lines in the tunnel TT2/TT2A has to be well coordinated.
Trigger for mercury system requested to be 24 V.
After lunch the the fiber cables are installed in order to allow checking the diagnostics. This check shall be done before the mercury fill. One fiber shows little transmission.
The (un-)mounting procedure of the fibers for transport does not seem to be well established. The fibers are exposed to a major mechanical handling. It could be considered to remove the fibers for transport to CERN to protect them from strong shocks. Consider putting the cable roll on the outside. At CERN it is required to open the secondary containment for adjusting and verifying the diagnostics.
It turned out, that the all fibers are in good shape. Mirrors are positioned by a system of springs and screws. During transport from ORNL to MIT the fine adjustment of the mirrors changed. This was fixed in the evening. Also the position of the mirrors was glued with epoxy.
1. March 2007 (Thursday)
A heater is placed on the outside of the secondary containment in order to replace the bore heater upon insertion of the snout into the bore. Unfortunately the heater is scratched off partly on the left side (beam view) during insertion and removal. Repair with some tape envisaged.
The cabling was changed inside the snout to protect the mirrors from being shifted.
The outer heat blanket will be replaced by two half pads leaving space at the place, where the heater was scratched off.
The afternoon was dedicated to the mercury filling.
Everything went well. Only vapor reading at putting the sensor close to the flask opening.
10 flasks poured corresponding to 21.45
Main items of filling procedure:
Active ventilation close to the handling.
Additional sheets and a pans prevent spill of mercury onto the floor.
Log book records the loaded mercury. Mercury flasks are weighted before and after filling.
No remarkable mercury vapor monitor reading.
2. March 2007 (Friday), IE and FH arrived
Temperatures are now stable above room temperature ins all locations of the snout inserted into the bore.
16:00 slow running in of mercury system.
18:00 each a 10, 15 and 20 m/s run, all fine, no solenoid pulsing yet.
3. March 2007 (Saturday)
The power supplies for the heaters showed troubles. Fuses replaced.
10:00: The area has been prepared for pulsing the solenoid (video: Overview of area)
Tests with a 20 m/s mercury jet and launching the emergency stop have been performed at ORNL. The procedure showed no troubles.
11:40 first shot with 10m/s and 3 Tesla (video: control room).
Conclusion: Running at 15 Tesla and 20 m/s successfully passed. The magnetic field clearly stabilizes the surface fluctuations. The few days showed, that the systems works as expected and handling could be done as foreseen. No safety problems occurred.
Tests completed. Removal scheduled on
Sunday and Monday.
4. March (Sunday)
Disconnecting the optical diagnostics
5. March 2007 (Monday)
The drainage of the mercury is performed similar to the filling.
Closing the snout (7 MB)
First flask transfer (17 MB)
Another flask transfer (6 MB)
Last flask and follow-up (12 MB)
Nominal test at 3 Tesla (2 MB)
Cryogenics leak (1 MB)